### v1

@@ 4.3 Symmetries with differential terms $Q_{(n)}$ (36) independent on spatial variables
Let us note that in fact we have to more equations additional to system (42). By definition function $M=\frac{1}{f}$ satisfies the condition (refer to (8))
@@
On the other hand, since function $g=g({\bf x})$ in (36) in fact is a function of the Euler angles otherwise the dilatation transformation will change such integral of motion. It follows from the above that function $N=gM$ satisfies the same additional condition as $M$, i.e., $x_{a}N_{a}=-2N$.
@@
The corresponding equation (42) is qualitatively different in two cases: $c_{1}c_{2}=0$ and $c_{1}c_{2}\neq 0$. In the first case we have only one nonzero coefficient, say, $c_{3}$. Setting $c_{3}=1$ we come to the following version of equation (42):
@@
$M=\frac{F(\varphi)}{\tilde{r}^{2}}+\frac{c_{1}}{x_{3}^{2}},\ N=-\frac{c_{1}}{x_{3}^{2}}$ (47)
where $F(\varphi)$ is an arbitrary function and $c_{1}$ is the integration constant.
Solutions of equations (21) are analogous to (47) but include another arbitrary function and integration constant. Substituting these solutions into formulae (22) we come to the inverse mass and potential presented in Item 1 of Table 2 where two integrals of motion are indicated, namely,
$Q=P_{3}^{2}-\left(\frac{c_{1}}{x_{3}^{2}}\cdot H\right)+\frac{c_{2}}{x_{3}^{2}}$ (48)
and
$\tilde{Q}=K_{3}^{2}-\left(\frac{c_{1}r^{4}}{x_{3}^{2}}\cdot H\right)+\frac{c_{2}r^{4}}{x_{3}^{2}}$ (49)
where we use the notation $(\Phi\cdot H)=P_{a}\Phi fP_{a}+\Phi V$. These integrals can be transformed one into another via the symmetry transformation (25).
In the case when both coefficients $c_{1}$ and $c_{2}$ are nontrivial equation (42) is reduced to the following system
$c_{1}M_{1}+N_{1}=0,\ c_{2}M_{3}+N_{3}=0,\ N_{2}=0$ (50)
@@
Comparing (51) with (47) it is not difficult to understand that in fact we have two symmetries corresponding to $c_{1}=0,c_{3}\neq 0$ and $c_{1}\neq 0,c_{3}=0$ and deal with their linear combination. Moreover, since all variables $x_{1},x_{2}$ and $x_{3}$ are involved into $M$ presented in (51) in the completely symmetric form we can predict one more symmetry whose differential part is equal to $P_{2}^{2}$. Just these results are presented in Item 5 of Table 2.
@@ 4.4 Symmetries with differential terms $Q_{(n)}$ (35) linear in spatial variables
$\mu^{ab}M_{b}+N_{a}=0$ (52)
since the first term in (19) appears to be identical zero.
By definition function $M$ and $N$ satisfy condition (43) and the following one:
@@
which can be used to solve equation (52) for $N$. Indeed, multiplying the latter equation by $x_{a}$, summing up with respect to the repeating index $a$ and using (43), (53) we obtain the following expression for $N$:
$N=r^{2}\lambda^{b}M_{b}+2\lambda^{b}x_{b}M$ (54)
or, using notations (8),
$N=\lambda^{b}F_{b}.$ (55)
Equations (52), (55) look rather gentle but in fact it is a rather complicated system including eight arbitrary parameters. Fortunately the number of these parameters can be reduced to five using the equivalence transformations which include rotations and dilatations. To do it let us mention that symmetry operator generated by the Killing tensor (38) can be represented as a bilinear combination of first order differential operators (24) in the following manner:
$Q=\lambda^{ab}\{P_{a},L_{b}\}+\lambda^{a}\{P_{a},D\}+\delta^{ab}g({\bf x}).$ (56)
Thus up to rotation transformations and the freedom in definition of their trace we can restrict ourselves to the following nonzero entries of tensor coefficients $\lambda^{ab}$:
$\lambda^{12}=\lambda^{21}=\mu,\ \lambda^{33}=\nu$ (57)
while the remaining entries can be reduced to zero.
The next reduction of the number of arbitrary parameters can be obtained using the following speculations. System (52) includes three linear equations for function $F$ dependent on two variables, and so to be consistent it should be degenerated. In other words, the determinant of the matrix whose entries are $\tilde{\mu}^{ab}$ should be equal to zero. This condition is satisfied in the following cases:
$\mu=0,\ \nu=0,\ \kappa=1,$ (58)
$\lambda_{1}=0,\ \lambda_{2}=0,\ \mu^{2}+\lambda_{3}^{2}-\nu^{2}=0,\ \kappa=0$ (59)
$\lambda_{2}=\pm\lambda_{1},\ \lambda_{3}=0,\ \nu^{2}\mp\nu\mu+\lambda_{1}^{2}=0,\ \kappa=0$ (60)
$\lambda_{1}=0,\ \lambda_{2}=0,\ \nu=0,\ \kappa=1,$ (61)
$\lambda_{1}=\pm\lambda_{2},\ \mu=\mp\nu,\ \lambda_{3}=0,\kappa=1$ (62)
while the parameters which are not included into (58)-(62) are arbitrary. In other words effectively we have four versions of system (54) any of which include as maximum two arbitrary parameters since up to rotation transformations in the case (58) we can restrict ourselves to the only nonzero parameter $\lambda^{3}$.
Let us consider versions (58)-(62) step by step. In the case (58) the only nonzero parameter is $\lambda^{3}$ and equation (52) is reduced to the following system:
@@
In addition, unknowns $M$ and $N$ should satisfy the following conditions:
$x_{a}M_{a}+2M=0,$ (64)
$x_{a}N_{a}+2N=0.$ (65)
System (63), (64) is easy solvable, its solutions are presented in Item 2 of Table 2, but in terms of functions (18).
Considering version (59) we find that the related equations (52) and (64) have nontrivial solutions only for the cases $\lambda_{3}=0,\mu=\pm\nu$ and $\lambda_{3}=\pm\nu,\mu=0$.
In the first case these equations are solved by the following functions:
$M=\frac{c_{1}}{(x_{1}\pm x_{2})^{2}+2x_{3}^{2}},N=0.$ (66)
The corresponding PDM Schrödinger equations admit two parametric Lie symmetry group including dilatations and simultaneous shifts of variables $x_{1}$ and $x_{2}$. Such systems are completely classified in [45].
In the second case we obtain the solution
$M=\frac{c_{1}}{x_{1}^{2}+x_{2}^{2}},\ N=0$ (67)
while for the case (69) we again have solutions (66) which also admits two Lie symmetries, i.e., dilatations and shifts along the third coordinate axis.
Version (60) generates a solvable system (52) and (64) in two cases:
$\lambda_{1}=0,\ \lambda_{2}=0,\ \lambda_{3}=0,\ \nu=0,\ \kappa=0,$ (68)
$\lambda_{1}=0,\ \lambda_{2}=0,\ \lambda_{3}=0,\ \mu=\pm\nu,\ \kappa=0$ (69)
In the case (68) the related system (52) and (64) is solved by functions (67) while for the case (69) we again have solutions (66). In other words we again have the systems with two parametric Lie symmetry groups.
The version (60) generates solutions (67). However, for the special cases $\lambda_{3}=\mu$ and $\lambda_{3}=0$ we obtain nice solutions presented in Items 3 of Table 1 and 2 of Table 2.
Table 1. Inverse masses, potentials and integrals of motion for integrable scale invariant systems.
No $f$ $V$ Integrals of motion
3 $\frac{\tilde{r}^{4}}{c_{1}(x_{3}^{2}+r^{2})e^{\pm 2\varphi}+c_{2}x_{3}\tilde{r}e^{\pm\varphi}+c_{3}\tilde{r}^{2}}$ $\frac{c_{4}(x_{3}^{2}+r^{2})e^{\pm 2\varphi}+c_{5}x_{3}\tilde{r}e^{\pm\varphi}+c_{6}\tilde{r}^{2}}{c_{1}(x_{3}^{2}+r^{2})e^{\pm 2\varphi}+c_{2}x_{3}\tilde{r}e^{\pm\varphi}+c_{3}\tilde{r}^{2}}$$\begin{array}[]{c}\{P_{3},(D\mp L_{3})\}\\ +2\left((\frac{c_{2}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{2}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}})\cdot H\right)\\ -2(\frac{c_{5}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{6}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}}),\\ \{K_{3},(D\mp L_{3})\}-15x_{3}\\ +2\left(r^{2}(\frac{c_{2}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{2}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}})\cdot H\right)\\ -2r^{2}(\frac{c_{5}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{6}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}})\end{array}$
In Table 1 and the following Tables 2–5 $F(.),G(.),\tilde{F}(.),\tilde{G}(.)$ are arbitrary functions whose arguments are fixed in the brackets, $\omega,z,\rho$ and $\Phi_{m}$ are given by relations (81) and (83), $c_{1},c_{2},...$ are arbitrary real parameters, $\hat{\eta}=3r^{2}+15x_{3}^{2}$.
4.5 Symmetries with differential terms $Q_{(n)}$ (36) quadratic in spatial variables

### v3

@@ 3.3 Symmetries with differential terms $Q_{(n)}$ (22) independent on spatial variables
Let us note that in fact we have two more equations additional to system (34). By definition function $M=\frac{1}{f}$ satisfies the condition (refer to (5))
@@
On the other hand, function $g=g({\bf x})$ in (22) should depend on the Euler angles and be independent on the radial variable otherwise the dilatation transformation will change such integral of motion. It follows from the above that function $N=gM$ satisfies the same additional condition as $M$, i.e.,
$x_{a}N_{a}=-2N.$ (36)
@@
The corresponding equation (34) is qualitatively different in two cases: $\mu\nu=0$ and $\mu\nu\neq 0$. In the first case we have only one nonzero coefficient, say, $\nu$. Setting $\nu=1$ we come to the following version of equation (34):
@@
$M=\frac{F(\varphi)}{\tilde{r}^{2}}-\frac{c_{1}}{x_{3}^{2}},\ N=\frac{c_{1}}{x_{3}^{2}}$ (40)
where $F(\varphi)$ is an arbitrary function, $\tilde{r}^{2}=x_{1}^{2}+x_{2}^{2}$ and $c_{1}$ is the integration constant.
Solutions of equations (26) are analogous to (40) but include another arbitrary function and integration constant. Substituting these solutions into formulae (23) we come to the inverse mass and potential presented in the following formulae
$f=\frac{x_{3}^{2}\tilde{r}^{2}}{F(\varphi)x_{3}^{2}-c_{1}\tilde{r}^{2}},\ V=\frac{G(\varphi)x_{3}^{2}+c_{2}\tilde{r}^{2}}{F(\varphi)x_{3}^{2}-c_{1}\tilde{r}^{2}}$ (41)
while the related integral of motion has the following form:
$Q=P_{3}^{2}-\left(\frac{c_{1}}{x_{3}^{2}}\cdot H\right)+\frac{c_{2}}{x_{3}^{2}}.$ (42)
where we use the notation
$(N\cdot H)=\sqrt{N}H\sqrt{N}$ (43)
(in our case $N=\frac{c_{1}}{x_{3}^{2}}$).
One more integral of motion can be obtained from (42) using symmetry transformation (28). It looks as follows:
$\tilde{Q}=K_{3}^{2}-\left(\frac{c_{1}r^{4}}{x_{3}^{2}}\cdot H\right)+\frac{c_{2}r^{4}}{x_{3}^{2}}+3r^{2}+15x_{3}^{2}$ (44)
In the case when both coefficients $\mu$ and $\nu$ are nontrivial equation (34) is reduced to the following system
$\begin{split}&\mu M_{1}+N_{1}=0,\\ &\nu M_{3}+N_{3}=0,\\ &N_{2}=0\end{split}$ (45)
@@
Comparing (46) with (40) it is not difficult to understand that in fact we have two symmetries corresponding to $c_{1}=0,c_{3}\neq 0$ and $c_{1}\neq 0,c_{3}=0$ and deal with their linear combination. Moreover, since all variables $x_{1},x_{2}$ and $x_{3}$ are involved into $M$ presented in (46) in the completely symmetric form we can predict one more symmetry whose differential part is equal to $P_{2}^{2}$.
Substituting solutions (46) into formulae (23) and (25) we obtain the following forms of the inverse mass and potential:
$\begin{split}&f=\frac{x_{1}^{2}x_{2}^{2}x_{3}^{2}}{c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2}},\\ &V=\frac{c_{4}x_{2}^{2}x_{3}^{2}+c_{5}x_{1}^{2}x_{3}^{2}+c_{6}x_{1}^{2}x_{2}^{2}}{c_{1}x_{2}^{2}x_{3}^{2}+c_{2}x_{1}^{2}x_{3}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}\end{split}$ (47)
The corresponding integrals of motion are similar to (42) and have the following form:
$\begin{split}&Q_{1}=P_{1}^{2}-\left(\frac{c_{1}}{x_{1}^{2}}\cdot H\right)+\frac{c_{4}}{x_{1}^{2}},\\ &Q_{2}=P_{2}^{2}-\left(\frac{c_{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{5}}{x_{2}^{2}},\\ &Q_{3}=P_{3}^{2}-\left(\frac{c_{3}}{x_{3}^{2}}\cdot H\right)+\frac{c_{6}}{x_{1}^{2}}.\end{split}$ (48)
Notice that only two integrals (48) are linearly independent since the condition $Q_{1}+Q_{2}+Q_{3}=0$ is satisfied. Two more linearly independent integrals of motion can be obtained by applying symmetry transformation (28) to $Q_{1}$ and $Q_{2}$.
@@ 3.4 Symmetries with differential terms $Q_{(n)}$ (21) linear in spatial variables
$R^{a}\equiv A^{ab}M_{b}+N_{a}=0$ (49)
where
$A^{ab}=(\mu^{ab}-\lambda^{a}x^{b}).$ (50)
We see that thanks to the presence of the additional condition generated by the dilatation symmetry equation (24) can be transformed to the linear and homogeneous algebraic constraint (50) for the first order derivatives $N_{a}$ an $M_{b}$.
3.4.1 Admissible combinations of arbitrary parameters
Derivatives $N_{a}$ present in equation (50) can be excluded. Moreover, it can be done in two independent ways. First, we can calculate the curl of vector $R^{a}$ and obtain the following second order system:
$(A^{ab}M_{b})_{c}=(A^{cb}M_{b})_{a}.$ (51)
One more way to exclude $N_{a}$ is to use the condition
@@ 3.4.1 Admissible combinations of arbitrary parameters
which should be satisfied to guarantee that the dilatation transformation does not destroy the intagral of motion. Indeed, multiplying equation (49) by $x_{a}$, summing up with respect to the repeating index $a$ and using (35), (52) we obtain the following expression for $N$:
$N=r^{2}\lambda^{b}M_{b}+\varepsilon^{bcd}\lambda^{cn}x_{n}x_{d}M_{b}.$ (53)
Substituting (53) into (49) we come to the following system for $M$:
$A^{ab}M_{b}+2x_{a}(\lambda^{b}M_{b})+r^{2}\lambda^{b}M_{ab}+\varepsilon^{bcd}\lambda^{cn}(x_{n}x_{d}M_{b})_{a}=0.$ (54)
Equations (35) , (49), (53) look rather gentle but in fact it is a rather complicated system including eight arbitrary parameters. Fortunately the number of these parameters can be reduced to five using the equivalence transformations which include rotations and dilatations. Indeed, like in previous section we can apply the rotations to reduce tensor components $\lambda^{ab}$ to the form presented in (38).
The next reduction of the number of arbitrary parameters can be obtained using the following speculations. System (51), (54) includes six linear equations for function $M$ dependent on two variables, refer to (5). Substituting the latter expression for $M$ into (51) and (54) we come to the system of six linear and homogeneous algebraic relations for variables $X_{1}=M_{\varphi},X_{2}=M_{\theta},X_{3}=M_{\varphi\varphi},X_{4}=M_{\varphi\theta},X_{5}=M_{\theta\theta}$ and $X_{6}M$ where $M_{\varphi}=\frac{\partial M}{\partial\varphi}$, etc., which has the following form:
$B^{\mu\nu}X_{\nu}=0,\mu,\nu=1,2,...,6,$ (55)
for the exact expressions of the matrix elements $B^{\mu\nu}$.
Equation (55) represents a necessary condition for the solvability of system (35) , (49) and (53). Thus to have a chance to obtain a nontrivial solution of the latter system we are supposed to choose such values of arbitrary parameters $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{11},\lambda_{11}$ which correspond to zero values of the determinant of the matrix whose entries are $B^{\mu\nu}$. The admissible sets of such parameters are given in the following formulae:
$\mu=0,\ \nu=0,\ \lambda_{a}{\text{ are arbitrary }},$ (56)
$\lambda_{1}=0,\ \lambda_{2}=0,\mu=0,\lambda_{3}=\pm\nu$ (57)
$\mu=-\nu,\lambda_{1}=0,\ \lambda_{2}=\pm\nu,\ \lambda_{3}=0$ (58)
$\lambda_{1}=\lambda_{2}=\lambda_{3}=0,\mu,\nu{\text{ are arbitrary }}.$ (59)
In other words effectively we have four versions of system (53) any of which include as maximum two arbitrary parameters since up to rotation transformations in the case (56) we can restrict ourselves to the only nonzero parameter $\lambda^{3}$.
3.4.2 Exact solutions for the arbitrary elements
Let us consider versions (56)-(58) step by step. In the case (56) we can choose the only nonzero parameter be $\lambda^{3}$ and equation (49) is reduced to the following system:
@@ 3.4.2 Exact solutions for the arbitrary elements
In addition, unknowns $M$ and $N$ should satisfy the conditions (35) and (52): which we rewrite once more for the readers convenience:
$\begin{split}&x_{a}M_{a}+2M=0,\\ &x_{a}N_{a}+N=0.\end{split}$ (61)
System (60), (61) is easy solvable, its solutions have the following form:
$M=\frac{F(\varphi)r+c_{1}x_{3}}{\tilde{r}^{2}r},\ N=\frac{c_{1}}{r}.$ (62)
while the corresponding integrals of motion look as:
$\begin{split}&Q_{1}=\{P_{3},D\}-\frac{c_{2}}{r}+\left(\frac{c_{1}}{r}\cdot H\right),\\ &Q_{2}=\{K_{3},D\}+(r\cdot H)-c_{2}r-15x_{3}\end{split}$ (63)
and are connected via the symmetry transformation (28).
Considering version (57) we come to the following form of equations (49):
$\begin{split}&\nu(\pm x_{1}-x_{2})M_{3}+N_{1}=0,\\ &\nu(x_{1}\pm x_{2})M_{3}+N_{2}=0,\end{split}$ (64)
$\nu(x_{1}M_{2}-x_{2}M_{1})\pm x_{3}M_{3}+N_{3}=0.$ (65)
The evident algebraic consequences of the subsystem (64) look as follows
$\begin{split}&\nu\tilde{r}^{2}M_{3}+x_{1}N_{1}+x_{2}N_{2}=0,\\ &\pm\nu\tilde{r}^{2}M_{3}+x_{1}N_{2}-x_{2}N_{1}=0.\end{split}$ (66)
In cylindric variables equations (66) and (65) take especially simple form:
$\begin{split}&\nu\tilde{r}^{2}M_{3}+\tilde{r}N_{\tilde{r}}=0,\\ &\pm\nu\tilde{r}^{2}M_{3}+N_{\varphi}=0,\\ &\nu M_{\varphi}\pm x_{3}M_{3}+N_{3}=0\end{split}$ (67)
where $M_{\tilde{r}}=\frac{\partial M}{\partial\tilde{r}}$, etc..
System (67) is completely equivalent to (64), (65). In addition, we have equations (61) whose form in the cylindric variables is:
$\begin{split}&\tilde{r}M_{\tilde{r}}+x_{3}M_{3}+2M=0,\\ &\tilde{r}N_{\tilde{r}}+x_{3}N_{3}+N=0.\end{split}$ (68)
The generic solutions of the overdetermined system (67), (78) have the following form:
$\begin{split}&M=\frac{c_{1}(x_{3}^{2}+r^{2})e^{\pm 2\varphi}+c_{2}x_{3}\tilde{r}e^{\pm\varphi}+c_{3}\tilde{r}^{2}}{\tilde{r}^{4}},\\ &N=\nu\left(\frac{c_{2}e^{\pm\varphi}}{\tilde{r}}+\frac{c_{1}x_{3}e^{\pm 2\varphi}}{\tilde{r}^{2}}\right).\end{split}$ (69)
The next step is to consider version (58). We choose the subversion $\lambda^{2}=+\nu$ which generates the following form of equations (49):
$\begin{split}&\nu((x_{1}+x_{3})M_{3}-2x_{2}M_{1})+N_{1}=0,\\ &\nu(x_{3}M_{1}+x_{2}M_{2}+x_{1}M_{3})+N_{2}=0,\\ &\nu(x_{1}+x_{3})M_{2}-2x_{2}M_{1}+N_{3}=0.\end{split}$ (70)
System (81) is solved by the following functions:
$\begin{split}&M=\frac{c_{1}}{y_{+}^{2}}+\frac{c_{2}}{y_{-}^{2}}+\frac{c_{3}x_{2}}{y_{+}^{2}\sqrt{y_{1}^{2}+2x_{2}^{2}}},\\ &N=\frac{c_{3}\sqrt{y_{+}^{2}+2x_{2}^{2}}}{y_{+}^{2}}+x_{2}\left(\frac{c_{1}}{y_{+}^{2}}+\frac{c_{2}}{y_{-}^{2}}\right)\end{split}$ (71)
where $y_{\pm}=x_{1}\pm x_{3}$.
Notice that the subversion $\lambda^{2}=-\nu$ generates the analogous solutions for $M$ and $N$ where, however, $y_{+}\to y_{-}$ and $y_{-}\to y_{+}$.
The next note is that applying the rotation around the second coordinate axis we can transform the pair $y_{+},y_{-}$ to $x_{1}$ and $x_{3}$. As a result we come to the following realization:
$\begin{split}&f=\frac{c_{1}}{x_{1}^{2}}+\frac{c_{2}}{x_{3}^{2}}+\frac{c_{3}x_{2}}{\tilde{r}x_{1}^{2}},\\ &N=\frac{c_{1}x_{2}}{x_{1}^{2}}+\frac{c_{3}(x_{1}^{2}+2x_{2}^{2})}{2\tilde{r}x_{1}^{2}}.\end{split}$ (72)
The remaining version (59) corresponds to the following form of equations (49):
$\begin{split}&x_{2}(\mu-\nu)M_{3}-\mu x_{3}M_{2}+N_{1}=0;\\ &\nu x_{1}M_{3}-\mu x_{3}M_{1}+N_{2}=0,\\ &x_{2}(\mu-\nu)M_{1}+\nu x_{1}M_{2}+N_{3}=0.\end{split}$ (73)
This system added by equations (61) has only trivial solutions provided $\mu$ and $\nu$ are arbitrary. For some special values of $\mu$ and $\nu$ there are nontrivial solutions which, however, are not interesting. Namely for $\mu=\nu,\mu=-\nu,\mu=0$ and $\mu=2\nu$ equations (84) are solved by
$M=\frac{c_{1}}{x_{2}^{2}+x_{3}^{2}},$
$M=\frac{c_{1}(x_{1}^{2}+x_{3}^{2})+c_{2}x_{1}x_{3}}{(x_{1}^{2}-x_{3}^{2})^{2}},$
$M=\frac{c_{1}}{x_{1}^{2}+x_{2}^{2}}$
and
$M=\frac{c_{1}(x_{1}^{2}+x_{2}^{2})+c_{2}x_{1}x_{2}}{(x_{1}^{2}-x_{2}^{2})^{2}}$
correspondingly.
Notice that any of the above presented functions $M$ depend on two variables. The related PDM systems are scale invariant and, in addition, are invariant w.r.t. shifts along some fixed lines and so admit two parametric Lie groups. Such systems are completely classified in [45] and we will not discuss them here.
Summarising, there are three inequivalent PDM systems which are scale invariant and admit second order integrals of motion generated by the Killing tensors linear in independent variables. The related arbitrary elements are given by formulae (62), (69) and (71). In addition, there are such systems admitting two parametric Lie groups which are enumerated in [45].
3.5 Symmetries with differential terms $Q_{(n)}$ (22) quadratic in independent variables